a) A Ford Mustang is selling for $30,000. What is the price in euros if the...

  1. a) A Ford Mustang is selling for $30,000. What is the price in euros if the exchange rate is 1.10 dollars per euro?

    b) A BMW is selling for 70,000 euros. What is the price in dollars if the exchange rate is 1.25 dollars per euro?
Answer #1

Answer

a)

Price in x currency =price in y currency * exchange rate in x currency

30000=Price in euro * 1.1

Price in euro =30000/1.1=27272.7273=27272.72 euros

=====

b)

Price in x currency =price in y currency * exchange rate in x currency

Price in dollars =70000*1.25=87500

the price in the dollar is $87500